# Focus on This, Not That! Steering LLMs With Adaptive Feature Specification

### v3

@@ Appendix F Additional Baselines Results
In addition to the baselines that we present in the main paper, Few-shot and $\text{SFT}(y_{\text{focus}})$, we include two additional baselines to further supplement these results. We give the complete list of baselines that we consider below:
@@
Zero-shot baseline. Finally, we include a zero-shot inference baseline using the original pre-trained models without additional fine-tuning on our spurious datasets. No in-context examples are used at inference time, and the model is not trained at all beyond it’s pre-training. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
@@
Few-shot baseline. This second baseline compares FIT training to few-shot inference using the original pre-trained models without additional fine-tuning on our spurious datasets. Specifically, we use 5 in-context examples across all datasets. For the in-context examples, we concatenate multiple examples one after the other, including the instructional prompt only for the first in-context example and the final test example. Each in-context example contains the same focus instruction as the test example for which they serve as context. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
@@
SFT($y_{\text{focus}}$) baseline. We implement an SFT baseline that follows the same training procedure as FIT, except during training, we exclude any focus instructions from the input prompts while still training on the focus labels. This provides a fair comparison with FIT, as the models are trained on the same input text and label pairs. The rest of the training setup, including hyperparameters and early stopping, remains identical to the FIT training setup. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
@@
SFT($y$) baseline. We implement a vanilla SFT baseline that simply trains a model using SFT on inputs and their ground truth labels (as opposed to focus labels in the SFT($y_{\text{focus}}$) baseline). During training, only standard IT prompts are used, with no additional focus instructions included. The rest of the training setup, including hyperparameters and early stopping, remains identical to the FIT training setup. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
@@
Figure 6: Full Baseline vs FIT Focus accuracy ($\uparrow$) on SS. Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the SS dataset.
Figure 7: Full Baseline vs FIT Focus accuracy ($\uparrow$) on SMNLI. Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the SMNLI dataset.
Figure 8: Full Baseline vs FIT Focus accuracy ($\uparrow$) on BBQ. Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the BBQ dataset.
Appendix G SMNLI Ablation of Training and Test time Focus Instruction Rephrasing Differences
We analyse the impact of using the same versus different sets of focus instructions at training and test time when applying FIT models. Specifically, we generate alternative test set focus instructions by paraphrasing the training focus instructions, as shown in Appendix E, using ChatGPT.
Figure 9: Focus accuracy ($\uparrow$) for different training and test focus instruction sets. Figure comparing focus accuracies ($\mathcal{A}_{\text{focus}}$) of sampling from the same (top) and different (bottom) sets of focus instructions at training and test time of models on the SMNLI dataset.
As depicted in Figure 9, the results of this ablation reveal negligible differences between using the same or different focus instruction phrasings during training and testing. This indicates that FIT effectively trains the model to focus on or ignore features, regardless of how the instructions are phrased.
Appendix H Comparison of FIT against a specific debiasing technique
FIT is a general framework designed to enable users to steer a model’s behavior based on specified features. This approach provides enhanced control over model outputs during inference, adding a critical layer of explainability and controllability to model predictions.
Results. The results of the debiasing experiment comparing FIT to the PoE method is shown in Figure 10. FIT performs equally as well as the PoE method as shown by the comparing the default prompt accuracy ($\emptyset$) for the PoE models against the $\text{focus}(C)$ results for the FIT models; both metrics correspond to causal accuracy for these prompt types. Indicating that FIT performs just as well as a dedicated debiasing technique.
However, the PoE method requires training two separate models and does not provide steerability at test time as shown by the low focus accuracy on $\text{focus}(S)$. Indeed the model defaults to the ground truth label across all prompt types and does not change behaviour despite different different focus specifications. This highlights the flexibility of FIT, which not only debiases effectively but also enables additional controllability during inference.
Figure 10: Focus Accuracy ($\uparrow$) of FIT against PoE Debiasing Technique. Figure showing the focus accuracies ($\mathcal{A}_{\text{focus}}$) of FIT (bottom row) and the dedicated debaiasing technique, PoE (top row), on the BBQ dataset.
Appendix I Spurious Sentiment (SS) Dataset
We take a pre-existing dataset, in this case SST-5 (Socher et al., 2013a), and modify it in order to induce a known spurious feature and create a spurious binary sentiment analysis dataset.
Data-generating process (DGP). We frame our DGP using a graphical model to describe the synthetic dataset that we create. We follow a similar model to that described in (Arjovsky et al., 2019), specifically the model used for generating their coloured MNIST dataset. We use the following variables within our graphical model:
@@ Appendix I Spurious Sentiment (SS) Dataset
  • $\tilde{S}$ proposed spurious feature sample, here this is the presence of the keywords Bayesian or Pineapple. We represent this as a binary vector $S\in\{0,1\}^{2}$, where the first and second components of this vector denote the presence of either the keyword Pineapple or Bayesian respectively. We restrict to consider only one keyword appearing in a given text at a time so that $\text{Val}(S)=\{(1,0),(0,1)\}$.
  • $S$ final spurious feature that is naturally inserted using a LLM into the final SS dataset example $X$. $S$ is a randomly flipped version of the proposed spurious feature $S$.
  • $\tilde{X}$ is a sampled example from the original dataset that we are modifying to inject known spurious correlations.
  • $X$ original example $\tilde{X}$ but augmented to include the spurious feature.
@@
Figure 11: SS DGP. Graphical model showing the data generating process for modifying examples from the SST-5 dataset to introduce a new spurious keyword feature $S$.
The graphical model describing the DGP of the SS dataset is given in Figure 11. This admits a functional representation in the form:
@@
$C\sim\texttt{Ber}(\rho_{C}),\text{ where }\rho_{C}=\rho_{C}(\mathcal{D});$ (17)
$\tilde{X}\sim p_{\mathcal{D}}(\cdot|C)\;,$ (18)
$\tilde{S}=\left(\bm{1}_{C=0},\bm{1}_{C=1}\right);$ (19)
$U_{S}\sim\texttt{Ber}(\rho_{\text{spurious}});$ (20)
$S=U_{S}\tilde{S}+(1-U_{S})(1-\tilde{S});$ (21)
$U_{\text{incld.}}\sim\texttt{Ber}(\rho_{\text{incld.}});$ (22)
$X=\texttt{LLM}(\tilde{X},S);$ (23)
@@
The variable $\rho_{C}$ gives the distribution of sentiment labels in the original binarised SST-5 dataset. Moreover, $p_{\mathcal{D}}(\tilde{x}|C)$ denotes the conditional dataset distribution of the different input texts give $C$ (here we assume that we are just uniformly sampling text with the given sentiment $C$) and $\bm{1}_{(\cdot)}$ denotes the indicator function.
Finally, we prove that $\rho_{\text{spurious}}$ gives the coocurrence rate/predictivity between the label $Y$ and the spurious feature $S$, and is well-defined notation in the sense that it corresponds to Definition 3.1 so that $\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s)$, where $y_{s}$ is the label that spurious feature value $s$ is spuriously correlated with.
Proposition I.1.
From the SCEs described above, assuming that we have a balanced label distribution, that is $\rho_{C}=1/2$, we have that
@@
First note that we have by Equation 17 and the assumption that $\rho_{C}=1/2$, that $\mathbb{P}(Y=1)=\mathbb{P}(Y=0)=1/2$. Moreover, using the partition theorem, we have that
$\mathbb{P}(S=s)$ $=\mathbb{P}(S=s\mid\tilde{S}=s)\mathbb{P}(\tilde{S}=s)+\mathbb{P}(S=s\mid\tilde{S}\neq s)\mathbb{P}(\tilde{S}\neq s)$ (26)
$=\rho_{\text{spurious}}\cdot\frac{1}{2}+(1-\rho_{\text{spurious}})\cdot\frac{1}{2}$ (27)
$=\frac{1}{2}.$ (28)
Finally note that $\mathbb{P}(S=s\mid Y=y_{s})=\rho_{\text{spurious}}$ as $Y=y_{s}$ forces $C=y_{s}$, and therefore that $\tilde{S}=s$ by construction.
Putting this all together and utilising Bayes’ rule gives
$=\frac{\mathbb{P}(S=s\mid Y=y_{s})\mathbb{P}(Y=y_{s})}{\mathbb{P}(S=s\mid\tilde{S}=s)\mathbb{P}(\tilde{S}=s)+\mathbb{P}(S=s\mid\tilde{S}\neq s)\mathbb{P}(\tilde{S}\neq s)}$ (30)
$=\frac{\rho_{\text{spurious}}\cdot\frac{1}{2}}{\frac{1}{2}}$ (31)
$=\rho_{\text{spurious}}.$ (32)
which gives the result. ∎
Within our experiments in Section 4.1, we always force the label distribution to be balanced, that is $\rho_{C}=1/2$, and assume that within each dataset split, $\rho_{\text{spurious}}$ is the same rate for all spurious feature values. Independence conditions during training for FIT. As specified in Appendix C, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features. Here, $Y_{S}$ is the spurious label spuriously correlated to spurious feature value $S$. The results below give sufficient conditions for these independence conditions to be specified with respect to the DGP described in Figure 11.
Proposition I.2.
Assuming the SCEs given above and the corresponding DGP described in Figure 11, if we have a uniform label distribution $p(Y=y)$, that is $\rho_{C}=1/2$, and have that $\rho_{\text{spurious}}=1/2$ in the SS training set, then we have that $Y\perp\!\!\!\perp S$.
@@
Proposition I.3.
Assuming the SCEs given above and the corresponding DGP described in Figure 11, if we have a uniform label distribution $p(Y=y)$, that is $\rho_{C}=1/2$, and have that $\rho_{\text{spurious}}=1/2$ in the SS training set, then we have that $Y_{S}\perp\!\!\!\perp C$.
Proof.
[…]
$\mathbb{P}(S=s\mid C=c)$ $=\sum_{\tilde{s}}\mathbb{P}(S=s\mid C=c,\tilde{S}=\tilde{s})\mathbb{P}(\tilde{S}=\tilde{s}\mid C=c)$ (33)
$=\sum_{\tilde{s}}\mathbb{P}(S=s\mid\tilde{S}=\tilde{s})\mathbb{P}(\tilde{S}=\tilde{s}\mid C=c)$ (34)
$=\underbrace{\mathbb{P}(S=s\mid\tilde{S}=s)}_{=\rho_{\text{spurious}}}\underbrace{\mathbb{P}(\tilde{S}=s\mid C=c)}_{=\frac{1}{2}}$ (35)
$+\underbrace{\mathbb{P}(S=s\mid\tilde{S}\neq s)}_{=1-\rho_{\text{spurious}}}\underbrace{\mathbb{P}(\tilde{S}\neq s\mid C=c)}_{\frac{1}{2}}$ (36)
$=\frac{1}{2}.$ (37)
[…]
SCM from this DGP. Through the above data generation process, we introduce a new spurious feature within the dataset $S$, the presence of the keywords Bayesian and Pineapple. Recalling that $S=(1,0)$ and $S=(0,1)$ correspond to insertion of the keywords Pineapple and Bayesian respectively, we introduce the following spurious correlations between feature values of $S$ and label $Y$:
  1. 1.
    The presence of the word Pineapple in the text $X$ is spuriously correlated the label 00 (negative sentiment).
  2. 2.
    The presence of the word Bayesian in the text $X$ is spuriously correlated with the label $1$ (positive sentiment).
The sentiment feature still remains core within the augmented SS dataset, fully predicting the label $Y$ for each dataset example.
The above DGP, through the introduction of spurious feature $S$, induces a SCM that describes the spurious correlation between spurious feature $S$ and the label $Y$. The SCM, shown in Figure 12, follows the style-content decomposition described in (Kaddour et al., 2022), where $U$ is some hidden confounding variable.
Data generation methodology. We use Llama-3.1-70B-Instruct to generate modifications $X$ of original dataset examples $\tilde{X}$ to create new text which include the new keywords feature. The prompt we use for generation when modifying examples to include spurious features is give as:
Figure 12: SS SCM. SCM showing showing the spurious correlation present between the keyword feature $S$ and the label $Y$ of examples within the SS dataset, induced through the described data augmentation process.
Appendix J Spurious NLI dataset (SMNLI)

### v4

@@ D.4 Complete Set of Baselines
In addition to the baselines that we present in the main paper, namely few-shot and $\text{SFT}(y_{\text{focus}})$, we include two additional baselines to further supplement these results. We give the complete list of baselines below:
@@
Zero-Shot Baseline. We include a zero-shot inference baseline using the original pre-trained models without additional fine-tuning on any dataset. No in-context examples are used at inference time. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
@@
Few-Shot Baseline. This second baseline compares FIT training to few-shot inference, using the original pre-trained models without additional fine-tuning on our spurious datasets. Specifically, we use 4 in-context examples across all datasets. For the in-context examples, we concatenate multiple examples one after the other. Each in-context example contains the same focus instruction as the test example for which they serve as context. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
@@
SFT($y_{\text{focus}}$) Baseline. We implement an SFT baseline that follows the same training procedure as FIT, except during training, we exclude any focus instructions from the input prompts while still training on the focus labels. This provides a fair comparison with FIT, as models are trained on the same input text-label pairs. The rest of the training setup, including hyperparameters and early stopping, remains identical to the FIT training setup. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
@@
SFT($y$) Baseline. We implement a vanilla SFT baseline that simply trains a model using SFT on inputs and their ground truth labels (as opposed to focus labels in the SFT($y_{\text{focus}}$) baseline). During training, only standard IT prompts are used corresponding to the default prompt $\emptyset$, with no additional focus instructions included. The rest of the training setup, including hyperparameters and early stopping, remains identical to the FIT training setup. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
@@
Appendix E Spurious Sentiment (SS) Dataset
We take a pre-existing dataset, in this case SST-5 (Socher et al., 2013a), and modify it in order to include a known spurious feature and create a spurious binary sentiment analysis dataset that we call the spurious sentiment (SS) dataset.
E.1 Data-generating process (DGP)
We frame our DGP using a graphical model to describe the synthetic dataset that we create. We follow a similar model to that described in (Arjovsky et al., 2019), specifically the model used for generating their coloured MNIST dataset. We use the following variables within our graphical model:
@@ E.1 Data-generating process (DGP)
  • $\tilde{S}$ proposed spurious feature sample, here this is the presence of the keywords “Pineapple” or “Bayesian”. We represent this as a binary categorical variable with $\text{Val}(\tilde{S})=\{\text{Pineapple},\text{Bayesian}\}$. We note that, this restricts us to consider only one keyword appearing in a text at any given time.
  • $S$ the final included spurious feature that is naturally inserted using a LLM into the final SS dataset example $X$. The feature $S$ is a randomly flipped version of the proposed spurious feature $\tilde{S}$. So here $\text{Val}(S)=\{\text{Pineapple},\text{Bayesian}\}$ also.
  • $\tilde{X}$ is a sampled example from the original SST-5 dataset.
  • $X$ original example $\tilde{X}$ but naturally augmented to include the spurious feature $S$, without changing the underlying sentiment of the example.
@@
The graphical model describing the DGP of the SS dataset is given in Figure 11(a). This admits a functional representation in the form:
@@
$C\sim\text{Ber}(\rho_{C}),\text{ where }\rho_{C}=\rho_{C}(\mathcal{D});$ (12)
$\tilde{X}\sim p_{\mathcal{D}}(\cdot|C)\;;$ (13)
$\tilde{S}=\begin{cases}\textit{Pineapple}&\text{if }C=0\\ \textit{Bayesian}&\text{if }C=1\end{cases};$ (14)
$U_{S}\sim\text{Ber}(\rho_{\text{spurious}});$ (15)
$S=\begin{cases}\tilde{S}&\text{if }U_{S}=1\\ \text{Val}(\tilde{S})\setminus\tilde{S}&\text{if }U_{S}=0\end{cases};$ (16)
$X=\text{LLM}(\tilde{X},S);$ (17)
@@
Here, Ber denotes a Bernoulli distribution, and the function LLM denotes a LLM that naturally injects a spurious feature $S$ into example $\tilde{X}$ without altering its underlying sentiment. The variable $\rho_{C}$ governs the distribution of sentiment labels in the original binarised SST-5 dataset. Moreover, $p_{\mathcal{D}}(\cdot|C)$ denotes the conditional dataset distribution of the different input texts given $C$ (here we assume that we are just uniformly sampling text with the given sentiment $C$).
(a) SS DGP. Graphical model showing the data generating process for modifying examples from the SS dataset to introduce a new spurious keyword feature, $S$.
(b) SS Causal Graph. Causal graph showing showing the spurious correlation present between the keyword feature $S$ and the label $Y$ for examples in the SS dataset, induced through the described data augmentation process.
Figure 11: SS DGP and Causal Graph. The DGP and associated causal graph describing the generation of dataset examples and showing the causal structure within examples.
Through the above DGP, we introduce a new spurious feature within the dataset, $S$. Recalling that $S=\text{Pineapple}$ and $S=\text{Bayesian}$ correspond to the insertion of the keywords Pineapple and Bayesian respectively, we introduce the following spurious correlations between the final included feature values of $S$ and label $Y$:
  1. (a)
    The presence of the word Pineapple in the text $X$, that is $S=\text{Pineapple}$, is spuriously correlated with the label 00 (negative sentiment). Therefore, the spurious label associated with $S=\text{Pineapple}$ is given as $y_{\text{Pineapple}}=0$.
  2. (b)
    The presence of the word Bayesian in the text $X$, that is $S=\text{Bayesian}$, is spuriously correlated with the label $1$ (positive sentiment). Therefore, the spurious label associated with $S=\text{Bayesian}$ is given as $y_{\text{Bayesian}}=1$.
The sentiment feature $C$ still remains core within the augmented SS dataset, fully predicting the label $Y$ for each dataset example.
Causal Graph from this DGP. The above DGP, through the introduction of spurious feature $S$, induces a causal graph that describes the spurious correlation between spurious feature $S$ and the label $Y$ in terms of the additional variables $X$ and $C$ only. The causal graph, shown in Figure 11(b), is similar to the style-content decomposition described in (Kaddour et al., 2022).
Showing that $\rho_{\text{spurious}}$ Corresponds to the Predictivity of $S$. We now prove that $\rho_{\text{spurious}}$ gives the co-occurrence rate/predictivity between the label $Y$ and the spurious feature $S$, and is well-defined notation in the sense that it corresponds to Definition 2 so that $\rho_{\text{spurious}}=\mathbb{P}(Y=y_{s}|S=s)$, where $y_{s}$ is the label that spurious feature value $s$ is spuriously correlated with. Note that this will hold constant irrespective of the feature value of $S$. We begin with the following proposition.
Proposition E.1.
From the DGP described above. we have that
$\mathbb{P}(Y=y_{s}\mid S=s)=\begin{cases}\displaystyle\frac{\rho_{C}\,\rho_{\text{spurious}}}{\rho_{C}\,\rho_{\text{spurious}}+(1-\rho_{C})\,(1-\rho_{\text{spurious}})}&\text{ if }s=\text{Bayesian},\\ \displaystyle\frac{(1-\rho_{C})\,\rho_{\text{spurious}}}{(1-\rho_{C})\,\rho_{\text{spurious}}+\rho_{C}\,(1-\rho_{\text{spurious}})}&\text{ if }s=\text{Pineapple}.\end{cases}$ (19)
Proof.
Using the partition theorem, we have that
$\mathbb{P}(S=s)$ $=\underbrace{\mathbb{P}(S=s\mid\tilde{S}=s)}_{=\mathbb{P}(U_{s}=1)}\mathbb{P}(\tilde{S}=s)+\underbrace{\mathbb{P}(S=s\mid\tilde{S}\neq s)}_{=\mathbb{P}(U_{s}=0)}\mathbb{P}(\tilde{S}\neq s)$ (20)
$=\mathbb{P}(U_{s}=1)\mathbb{P}(\tilde{S}=s)+\mathbb{P}(U_{s}=0)\mathbb{P}(\tilde{S}\neq s)$ (21)
$=\rho_{\text{spurious}}\cdot\mathbb{P}(\tilde{S}=s)+(1-\rho_{\text{spurious}})\cdot\mathbb{P}(\tilde{S}\neq s)$ (22)
=$=\begin{cases}\rho_{C}\,\rho_{\text{spurious}}+(1-\rho_{C})\,(1-\rho_{\text{spurious}})&\text{ if }s=\textit{Bayesian},\\ (1-\rho_{C})\,\rho_{\text{spurious}}+\rho_{C}\,(1-\rho_{\text{spurious}})&\text{ if }s=\textit{Pineapple}.\end{cases},$ (23)
where we have used that $\mathbb{P}(U_{s}=1)=\rho_{\text{spurious}}$, and that the value of $\tilde{S}$ depends solely on the value of $C$.
Using this alongside Bayes’ rule gives
$=\frac{\mathbb{P}(S=s\mid C=y_{s})\mathbb{P}(Y=y_{s})}{\mathbb{P}(S=s)}$ (25)
$=\frac{\mathbb{P}(S=s\mid\tilde{S}=s)\mathbb{P}(Y=y_{s})}{\mathbb{P}(S=s)}$ (26)
=$=\begin{cases}\displaystyle\frac{\rho_{C}\,\rho_{\text{spurious}}}{\rho_{C}\,\rho_{\text{spurious}}+(1-\rho_{C})\,(1-\rho_{\text{spurious}})}&\text{ if }s=\text{Bayesian},\\ \displaystyle\frac{(1-\rho_{C})\,\rho_{\text{spurious}}}{(1-\rho_{C})\,\rho_{\text{spurious}}+\rho_{C}\,(1-\rho_{\text{spurious}})}&\text{ if }s=\text{Pineapple}.\end{cases}$ (27)
which gives the result. ∎
Corollary E.1.
From the DGP described above, assuming that we have a balanced label distribution, that is $\rho_{C}=1/2$, we have that
@@
This follows immediately from he previous proposition, Proposition E.1, with $\rho_{C}=1/2$. ∎
Within our experiments on the SS datsaet, we always force the label distribution to be balanced, that is $\rho_{C}=1/2$, and assume that within each dataset split, $\rho_{\text{spurious}}$ is the same rate for all spurious feature values.
Data Generation Methodology. We use Llama-3.1-70B-Instruct to generate modifications $X$ of original dataset examples $\tilde{X}$ to create new text which include the new keyword features (presence of the keywords “Bayesian” and “Pineapple”. The prompt we use for generation when modifying examples to include spurious features is give as:
E.2 Independence Conditions During Training for FIT on SS.
As specified in Appendix D, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features. Here, $Y_{S}$ is the spurious label spuriously correlated to the random spurious feature value $S$. The results below give sufficient conditions for these independence conditions to be satisfied with respect to the DGP described above, and consequently form the conditions that we impose on the SS training set for FIT training. We show that the key condition that we must impose in training for $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ is to set $\rho_{\text{spurious}}=1/2$.
Proposition E.2.
Assuming the DGP described in above and assuming that $\rho_{\text{spurious}}=1/2$, we have that $Y\perp\!\!\!\perp S$.
Proof.
With $\rho_{\text{spurious}}=1/2$, we have from the DGP given above, that $\tilde{S}\perp\!\!\!\perp S$. In particular, using the factorisation implied by the directed acyclic graph (DAG) corresponding to the DGP, we see that
$\mathbb{P}(Y=y,S=s)$ $=\sum_{c\in\text{Val(C)},\tilde{s}\in\text{Val}(\tilde{S})}\mathbb{P}(Y=y\mid C=c)\mathbb{P}(C=c)\mathbb{P}(\tilde{S}=\tilde{s}\mid C=c)\underbrace{\mathbb{P}(S=s\mid\tilde{S}=\tilde{s})}_{=\mathbb{P}(S=s)\text{ as}S\perp\!\!\!\perp\tilde{S}}$ (29)
$=\sum_{c\in\text{Val(C)},\tilde{s}\in\text{Val}(\tilde{S})}\mathbb{P}(Y=y\mid C=c)\mathbb{P}(C=c)\mathbb{P}(\tilde{S}=\tilde{s}\mid C=c)\mathbb{P}(S=s)$ (30)
$=\mathbb{P}(S=s)\sum_{c\in\text{Val(C)},\tilde{s}\in\text{Val}(\tilde{S})}\mathbb{P}(Y=y\mid C=c)\mathbb{P}(C=c)\mathbb{P}(\tilde{S}=\tilde{s}\mid C=c)$ (31)
$=\mathbb{P}(S=s)\sum_{c\in\text{Val(C)}}\mathbb{P}(Y=y\mid C=c)\mathbb{P}(C=c)$ (32)
$=\mathbb{P}(S=s)\mathbb{P}(Y=y).$ (33)
By definition, this shows that we have that $Y\perp\!\!\!\perp S$, as required. ∎
Proposition E.3.
Assuming the DGP described above, for $\rho_{\text{spurious}}=1/2$, then we have that $Y_{S}\perp\!\!\!\perp C$.
@@
E.3 Results
Figure 12 (a) shows the focus accuracy results of three LLMs on the SS dataset across all of the baseline methods described in Section D.4 and the FIT method. We see that across all focus instructions and all models, FIT shows significant improvement over the baselines, achieving very high focus accuracy across all focus instruction types and across all test sets with varying predictivity levels.
Table 4: Complete Baselines vs. FIT Focus Accuracies on SS ($\uparrow$). For each method, we report the mean of focus‐accuracy means ($\mathcal{A}_{\text{focus}}$) over the four test splits ($\mathcal{D}_{\text{iid}}$, $\mathcal{D}_{\text{high}}$, $\mathcal{D}_{\text{low}}$ and $\mathcal{D}_{\text{flipped}}$) $\pm$ the between‐split standard deviation of these means (i.e. how performance varies as predictivity changes) across repeats. Boldface indicates the best mean for each focus instruction type and model independently.
$\varnothing$ $\mathrm{focus}(C)$ $\mathrm{focus}(C)\wedge\mathrm{ignore}(S)$ $\mathrm{ignore}(S)$ $\mathrm{focus}(S)$ $\mathrm{focus}(S)\wedge\mathrm{ignore}(C)$
Mistral Zero-shot $0.886_{\pm 0.007}$ $0.909_{\pm 0.004}$ $0.899_{\pm 0.002}$ $0.871_{\pm 0.006}$ $0.423_{\pm 0.199}$ $0.305_{\pm 0.104}$
Few-shot $0.893_{\pm 0.003}$ $0.906_{\pm 0.005}$ $0.904_{\pm 0.006}$ $0.793_{\pm 0.036}$ $0.559_{\pm 0.168}$ $0.634_{\pm 0.116}$
SFT($y$) $0.951_{\pm 0.005}$ $0.950_{\pm 0.004}$ $0.952_{\pm 0.005}$ $0.951_{\pm 0.004}$ $0.445_{\pm 0.275}$ $0.447_{\pm 0.270}$
SFT($y_{\mathrm{focus}}$) $0.751_{\pm 0.126}$ $0.794_{\pm 0.108}$ $0.903_{\pm 0.042}$ $0.879_{\pm 0.054}$ $0.506_{\pm 0.248}$ $0.519_{\pm 0.241}$
FIT $\mathbf{0.953_{\pm 0.004}}$ $\mathbf{0.954_{\pm 0.004}}$ $\mathbf{0.955_{\pm 0.004}}$ $\mathbf{0.954_{\pm 0.004}}$ $\mathbf{0.999_{\pm 0.001}}$ $\mathbf{0.999_{\pm 0.001}}$
Llama Zero-shot $0.500_{\pm 0.000}$ $0.500_{\pm 0.000}$ $0.500_{\pm 0.000}$ $0.500_{\pm 0.000}$ $0.506_{\pm 0.003}$ $0.506_{\pm 0.002}$
Few-shot $0.674_{\pm 0.006}$ $0.838_{\pm 0.008}$ $0.630_{\pm 0.018}$ $0.508_{\pm 0.005}$ $0.491_{\pm 0.047}$ $0.502_{\pm 0.038}$
SFT($y$) $\mathbf{0.952_{\pm 0.004}}$ $\mathbf{0.954_{\pm 0.003}}$ $\mathbf{0.952_{\pm 0.007}}$ $\mathbf{0.952_{\pm 0.008}}$ $0.445_{\pm 0.276}$ $0.444_{\pm 0.272}$
SFT($y_{\mathrm{focus}}$) $0.668_{\pm 0.178}$ $0.686_{\pm 0.166}$ $0.803_{\pm 0.095}$ $0.795_{\pm 0.094}$ $0.606_{\pm 0.197}$ $0.601_{\pm 0.193}$
FIT $0.949_{\pm 0.002}$ $0.951_{\pm 0.002}$ $0.952_{\pm 0.002}$ $0.950_{\pm 0.002}$ $\mathbf{0.998_{\pm 0.001}}$ $\mathbf{0.999_{\pm 0.001}}$
Vicuna Zero-shot $0.420_{\pm 0.012}$ $0.584_{\pm 0.007}$ $0.381_{\pm 0.008}$ $0.355_{\pm 0.006}$ $0.199_{\pm 0.105}$ $0.129_{\pm 0.066}$
Few-shot $0.147_{\pm 0.010}$ $0.459_{\pm 0.014}$ $0.533_{\pm 0.012}$ $0.431_{\pm 0.010}$ $0.300_{\pm 0.150}$ $0.300_{\pm 0.125}$
SFT($y$) $\mathbf{0.955_{\pm 0.005}}$ $\mathbf{0.956_{\pm 0.005}}$ $0.955_{\pm 0.004}$ $0.953_{\pm 0.006}$ $0.446_{\pm 0.278}$ $0.445_{\pm 0.277}$
SFT($y_{\mathrm{focus}}$) $0.570_{\pm 0.180}$ $0.602_{\pm 0.187}$ $0.685_{\pm 0.138}$ $0.671_{\pm 0.129}$ $0.676_{\pm 0.129}$ $0.660_{\pm 0.118}$
FIT $0.950_{\pm 0.003}$ $0.953_{\pm 0.003}$ $\mathbf{0.955_{\pm 0.003}}$ $\mathbf{0.955_{\pm 0.004}}$ $\mathbf{0.999_{\pm 0.001}}$ $\mathbf{0.999_{\pm 0.001}}$
Figure 12: Complete Baseline vs FIT Focus Accuracies on SS ($\uparrow$). Figure giving the mean focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the SS dataset. The maximum standard deviations of FIT, $\text{SFT}(y_{\text{focus}})$, $\text{SFT}(y)$ and few-shot methods across models are 2.17%, 14.8%, 0.65%, and 2.83% respectively. fcs = focus, ign = ignore.
Key Takeaways. High focus accuracy on SS indicates that FIT successfully steers model responses based on the feature which it is instructed to focus or to not focus on.
Appendix F Spurious NLI dataset (SMNLI)